When 10.0 mg of sucrose is dissolved in enough water to make 100.0 mL of solution at 25.00 °C, an osmotic pressure of 5.43 mmHg is measured. What is the molar mass of sucrose? Show your work, and provide your answer in g/mol.
Answer #1

P= 5.43 mm Hg

= (5.43/760) atm

= 0.0071 atm

T= 25.0 oC

= (25.0+273) K

= 298 K

use:

P = C*R*T

7.145*10^-3 = C*0.08206*298.0

C =2.922*10^-4 M

volume , V = 1*10^2 mL

= 0.1 L

use:

number of mol,

n = Molarity * Volume

= 2.922*10^-4*0.1

= 2.922*10^-5 mol

mass(solute)= 10.0 mg

= 0.01 g

use:

number of mol = mass / molar mass

2.922*10^-5 mol = (0.01 g)/molar mass

molar mass = 3.423*10^2 g/mol

Answer: 3.42*10^2 g/mol